# An Evolutionary Algorithm to Accelerate High-Dimensional Optimization

### v3

@@ Abstract.
Evaluated on the notoriously difficult CEC2017 benchmark suite of 29 test functions, QUASAR achieved the lowest overall rank sum (367) using the Friedman test, outperforming DE (735) and L-SHADE (452). Geometric mean comparisons show average final solution quality improvements of $3.85\times$ and $2.07\times$ compared to DE and L-SHADE, respectively ($p\ll 0.001$), with average optimization speed averaging $1.40\times$ and $5.16\times$ faster.
QUASAR’s performance establishes it as an effective, efficient, and user-friendly evolutionary algorithm for complex high-dimensional problems.
copyright: none
1. Introduction
High-dimensional numerical optimization remains a persistent challenge across many fields, including machine learning, science, engineering, and finance (Zhou et al., 2024). This challenge stems from the curse of dimensionality; the exponential growth of the search space renders gradient-based and parametric search methods ineffective (Shan and Wang, 2008). Consequently, optimization metaheuristics, particularly evolutionary algorithms, have become essential for locating optimal solutions in complex function landscapes.
Among these methods, Differential Evolution (DE) has established itself as a robust and highly effective optimizer due to its simplicity, efficiency, and effective mutation strategies (Storn, 1996). However, standard DE variants can struggle to maintain an effective balance between global exploration and local exploitation, particularly in high-dimensional search spaces. This often leads to stagnation or premature convergence that requires frequent restarts or extensive hyperparameter tuning, a significant practical drawback for real-world applications.
@@ 1. Introduction
The efficacy of QUASAR is demonstrated through rigorous testing against standard DE and its state-of-the-art variant L-SHADE across the complete CEC2017 benchmark suite. The results show large and statistically significant performance gains; these improvements are achieved using default parameters, highlighting QUASAR’s ease of use.
The remainder of this paper is organized as follows: Section 2 details the QUASAR algorithm, including its mutation, crossover, selection, and reinitialization mechanisms. Section 3 defines the experimental design, and Section 4 analyzes the experimental results compared to DE and L-SHADE. Section 5 describes the ease of use and implementation of the algorithm. Section 6 concludes the paper, with Section 7 proposing future work.
2. QUASAR Algorithm
@@
2.1. Initial Population
The initial population is generated as a Sobol sequence by default, due to its low discrepancy and superior uniformity in high-dimensional spaces compared to other quasi-Monte Carlo sampling methods (Bonneel et al., 2025). The default population size, $N$, is set to $10\times D$, where $D$ is the number of dimensions.
Additional initialization methods include uniform Latin Hypercube and random distributions, as well as non-uniform Hyperellipsoid Density sequences (Soltes, 2025).
2.2. Mutation
@@
2.2.1. Local Mutation (Exploitation):
@@
(Spooky-Best) $\mathbf{v}_{i,g}=\mathbf{X}_{\text{best},g}+F_{\text{local}}(\mathbf{X}_{i,g}-\mathbf{X}_{\text{rand},g})$
@@
2.2.2. Global Mutations (Exploration):
@@
(Spooky-Current) $\mathbf{v}_{i,g}=\mathbf{X}_{i,g}+F_{\text{global}}(\mathbf{X}_{\text{best},g}-\mathbf{X}_{\text{rand},g})$
(Spooky-Random) $\mathbf{v}_{i,g}=\mathbf{X}_{\text{rand},g}+F_{\text{global}}(\mathbf{X}_{i,g}-\mathbf{X}_{\text{rand},g})$
@@
Figure 1. Mutation factor distributions for global and local mutation strategies.
2.3. Crossover
QUASAR uses binomial crossover with a dynamic crossover rate ($CR_{i}$) proportional to the solution’s fitness rank. The best solutions are rapidly recombined to maximize both exploration and exploitation via synergy with global and local mutations.
@@ 2.3. Crossover
(1) $CR_{\text{raw},i}=\frac{N-1-\text{rank}_{i}}{N-1}$
(2) $CR_{i}=\max(CR_{\text{raw},i},0.33)$
@@
(3) $u_{i,n}=\begin{cases}v_{i,n},&\text{if }\text{rand}(0,1)\leq CR_{i}\\ X_{i,n},&\text{if }\text{rand}(0,1)>CR_{i}\end{cases}$
2.4. Selection
@@
(4) $\mathbf{X}_{i,(g+1)}=\begin{cases}\mathbf{u}_{i},&\text{if }f(\mathbf{u}_{i})<f(\mathbf{X}_{i})\\ \mathbf{X}_{i},&\text{if }f(\mathbf{u}_{i})\geq f(\mathbf{X}_{i})\end{cases}$
2.5. Asymptotic Reinitialization
@@
(5) $P_{\text{reinit}}(g)=\exp\left(\frac{\ln(P_{\text{final}})}{g_{\text{final}}\cdot g_{\text{max}}}\cdot g\right)$
@@
(6) $\boldsymbol{\mu}=\frac{1}{M}\sum_{i=1}^{M}\mathbf{X}_{i}$
(7) $\boldsymbol{\Sigma}=\frac{1}{M-1}\sum_{i=1}^{M}(\mathbf{X}_{i}-\boldsymbol{\mu})(\mathbf{X}_{i}-\boldsymbol{\mu})^{\top}+\epsilon\mathbf{I}$
@@
(8) $\mathbf{y}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})+\boldsymbol{\delta}$
@@
(9) $\boldsymbol{\delta}\sim\mathcal{N}\left(0,\left(\frac{\mathbf{b}_{\text{high},n}-\mathbf{b}_{\text{low},n}}{20}\right)^{2}\mathbf{I}\right)$
@@
(10) $\mathbf{X}_{\text{new},n}=\text{clip}(\mathbf{y}_{n},\mathbf{B}_{n})$
@@
Figure 2. QUASAR workflow for evolving each generation.
3. Experimental Design
3.1. Optimizers & Test Functions
@@
This choice is justified by the fact that the QUASAR algorithm is built on the DE framework, as well as DE’s standing as a robust and effective optimizer (Price, 1996). L-SHADE was specifically chosen due to its success in CEC optimization competitions (Mo et al., 2019).
@@
3.2. Experimental Setup
To ensure a fair comparison, each optimizer was run using their default hyperparameters. The initial population generation methods were also left to defaults; the tested sample sizes not being powers of 2 diminishes the uniformity of QUASAR’s initial Sobol sequence (Joe and Kuo, 2008), providing a theoretical initial handicap.
@@ 3.2. Experimental Setup
The computational budget for metaheuristic comparison is often defined by the maximum number of objective function evaluations ($FES_{max}$) (Kazikova et al., 2021). Due to L-SHADE’s linear population reduction, the $FES_{max}$ becomes a difficult comparison metric to compare against. As such, the primary comparisons are performed on final solution fitness errors and overall optimization times, given the same number of generations, using modern implementations via SciPy’s DE and MealPy’s L-SHADE.
@@ 3.3. Experimental Trials:
3.3.1. Dimension-Variant Trials:
The dimension-variant trials assessed QUASAR’s scalability by scaling the problem dimension $D$ with a fixed population size ($N=1000$). $30$ independent trials were performed for each test function/dimension combination to ensure statistical significance.
@@ 3.3.1. Dimension-Variant Trials:
3.3.2. Sample-Variant Trials:
The sample-variant trials assessed the impact of population size on performance. The computational budget is varied between population sizes of $[100,250,500,1000,5000]$. $30$ independent trials were performed for each test function / sample size combination, using maximum number of iterations $g_{max}=100$ and fixed dimensionality $D=100$.
@@ 3.3.2. Sample-Variant Trials:
Table 1. Experimental Parameters for Sample and Dimension Variations
@@
3.4. Statistical Analysis:
3.4.1. Overall Rank:
@@
3.4.2. Solution Quality Improvements:
@@ 3.4.3. Computational Efficiency:
(11) $\bar{R}{\text{time}}=\sqrt[k]{\prod{j=1}^{k}R_{\text{time},j}}$
3.4.4. Statistical Significance:
@@ 4. Results
Table 2. Rank Sums by Sample Size and Dimension
@@
Table 3. Overall Performance Comparison (Geometric Mean Ratios)
@@ 4.1.1. Dimension-Variant:
Figure 3. Geometric mean error (log-scale) by dimension for each optimizer, with population size of 1000.
4.1.2. Sample-Variant:
To assess the impact of varying population sizes, QUASAR and DE were compared across a fixed $100D$ and 100 maximum generations $g_{max}$, using various sample sizes ($N$).
@@ 4.1.2. Sample-Variant:
Figure 4. Geometric mean error (log-scale) by population size for each optimizer, in 100 dimensions.
Table 4. Solution Quality Comparison (Geometric Mean Error)
@@ 4.2.1. Dimension-Variant:
Figure 5. Average optimization run times for each optimizer, by dimension.
4.2.2. Sample-Variant:
@@
Figure 6. Average optimization run times for each optimizer, by population size.
Table 5. Efficiency Comparison (Geometric Mean Run Time)
@@
5. Ease of Use and Implementation
5.1. Ease of Use
@@ 5.2. Implementation
The source code requires minimal dependencies, relying only on NumPy and SciPy. The open-source package containing QUASAR, hdim_opt, is available via PyPi or at https://github.com/jgsoltes/hdim-opt, which includes the experimental notebooks for full reproducibility.
6. Conclusion
@@
7. Future Works
7.1. Enhancements
QUASAR’s quasi-adaptive mechanisms are highly effective; future improvements will explore fully self-adaptive strategies, similar to those found in other modern DE variants.
7.1.1. Mutations:
@@ 7.1.4. Verification:
Acknowledgements.
The author would like to thank Dr. Kellen Sorauf, Dr. Ksenia Polson, Dr. Mike Busch, and Monet Morris for their invaluable feedback and support throughout this study. It would not have been possible without the insights gained from each of our discussions. Dr. Sorauf’s recommendation to include L-SHADE in the experiment significantly improved the rigor of the study; Dr. Polson’s lectures laid the framework for the statistical analysis; Dr. Busch pushed me to pursue the project to begin with. Monet remains a source of inspiration, as always. This research was conducted independently and was not part of a formal project.

### v4

# An Evolutionary Algorithm to Accelerate High-Dimensional Numerical Optimization

@@ Abstract
Evaluated on the notoriously difficult CEC2017 benchmark suite of 29 test functions, QUASAR achieved the lowest overall rank sum (367) using the Friedman test, outperforming DE (735) and L-SHADE (452). Geometric mean comparisons show average final solution quality improvements of $3.85\times$ and $2.07\times$ compared to DE and L-SHADE, respectively ($p\ll 0.001$), with average optimization speed averaging $1.40\times$ and $5.16\times$ faster. QUASAR’s performance establishes it as an effective, efficient, and user-friendly evolutionary algorithm for complex high-dimensional problems.
keywords:
Evolutionary computation , Metaheuristics , Global optimization , Quantum-inspired algorithm , Covariance reinitialization , High-dimensional optimization
journal: Swarm and Evolutionary Computation
[…]
organization=Regis University,addressline=3333 Regis Blvd., city=Denver, postcode=80221, state=Colorado, country=USA
1 Introduction
High-dimensional numerical optimization remains a persistent challenge across many fields, including machine learning, science, engineering, and finance [9]. This challenge stems from the curse of dimensionality; the exponential growth of the search space renders gradient-based and parametric search methods ineffective [6]. Consequently, optimization metaheuristics, particularly evolutionary algorithms, have become essential for locating optimal solutions in complex function landscapes.
Among these methods, Differential Evolution (DE) has established itself as a robust and highly effective optimizer due to its simplicity, efficiency, and effective mutation strategies [8]. However, standard DE variants can struggle to maintain an effective balance between global exploration and local exploitation, particularly in high-dimensional search spaces. This often leads to stagnation or premature convergence that requires frequent restarts or extensive hyperparameter tuning, a significant practical drawback for real-world applications.
@@ 1 Introduction
The efficacy of QUASAR is demonstrated through rigorous evaluation against standard DE and its state-of-the-art variant L-SHADE across the full CEC2017 test function benchmark suite. The results show large and statistically significant improvements in final solution quality, with high computational efficiency. These performance gains are seen using its default parameters, highlighting the ease of use and scalability of the algorithm.
The remainder of the paper is structured as follows: Section 2 details the QUASAR algorithm, including its mutation, crossover, selection, and reinitialization mechanisms. Section 3 defines the experimental design and statistical analysis, Section 4 analyzes the results. Section 5 describes the ease of use and implementation of the algorithm. Section 6 summarizes the results and concludes the paper, with Section 7 proposing future enhancements and validations.
2 QUASAR Algorithm
@@
2.1 Initial Population
The initial population is generated as a Sobol sequence by default, due to its low discrepancy and superior uniformity in high-dimensional spaces compared to other quasi-Monte Carlo sampling methods [1]. The default population size, $N$, is set to $10\times D$, where $D$ is the number of dimensions.
Additional initialization methods include uniform Latin Hypercube and random distributions, as well as non-uniform Hyperellipsoid Density sequences [7].
2.2 Mutation
@@
2.2.1 Local Mutation (Exploitation):
@@
$\mathbf{v}_{i,g}=\mathbf{X}_{\text{best},g}+F_{\text{local}}(\mathbf{X}_{i,g}-\mathbf{X}_{\text{rand},g})$ (Spooky-Best)
@@
2.2.2 Global Mutations (Exploration):
@@
$\mathbf{v}_{i,g}=\mathbf{X}_{i,g}+F_{\text{global}}(\mathbf{X}_{\text{best},g}-\mathbf{X}_{\text{rand},g})$ (Spooky-Current)
$\mathbf{v}_{i,g}=\mathbf{X}_{\text{rand},g}+F_{\text{global}}(\mathbf{X}_{i,g}-\mathbf{X}_{\text{rand},g})$ (Spooky-Random)
@@
Figure 1: Mutation factor distributions for global and local mutation strategies.
2.3 Crossover
QUASAR uses binomial crossover with a dynamic crossover rate ($CR_{i}$) inversely proportional to the solution’s fitness rank. The best solutions are rapidly recombined to maximize both exploration and exploitation through synergy with global and local mutations.
@@ 2.3 Crossover
$CR_{\text{raw},i}=\frac{N-1-\text{rank}_{i}}{N-1}$ (1)
$CR_{i}=\max(CR_{\text{raw},i},0.33)$ (2)
@@
$u_{i,n}=\begin{cases}v_{i,n},&\text{if }\text{rand}(0,1)\leq CR_{i}\\ X_{i,n},&\text{if }\text{rand}(0,1)>CR_{i}\end{cases}$ (3)
2.4 Selection
@@
$\mathbf{X}_{i,(g+1)}=\begin{cases}\mathbf{u}_{i},&\text{if }f(\mathbf{u}_{i})<f(\mathbf{X}_{i})\\ \mathbf{X}_{i},&\text{if }f(\mathbf{u}_{i})\geq f(\mathbf{X}_{i})\end{cases}$ (4)
2.5 Asymptotic Reinitialization
@@
$P_{\text{reinit}}(g)=\exp\left(\frac{\ln(P_{\text{final}})}{g_{\text{final}}\cdot g_{\text{max}}}\cdot g\right)$ (5)
@@
$\boldsymbol{\mu}=\frac{1}{M}\sum_{i=1}^{M}\mathbf{X}_{i}$ (6)
$\boldsymbol{\Sigma}=\frac{1}{M-1}\sum_{i=1}^{M}(\mathbf{X}_{i}-\boldsymbol{\mu})(\mathbf{X}_{i}-\boldsymbol{\mu})^{\top}+\epsilon\mathbf{I}$ (7)
@@
$\mathbf{y}\sim\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})+\boldsymbol{\delta}$ (8)
@@
$\boldsymbol{\delta}\sim\mathcal{N}\left(0,\left(\frac{\mathbf{b}_{\text{high},n}-\mathbf{b}_{\text{low},n}}{20}\right)^{2}\mathbf{I}\right)$ (9)
@@
$\mathbf{X}_{\text{new},n}=\text{clip}(\mathbf{y}_{n},\mathbf{B}_{n})$ (10)
@@
Figure 2: QUASAR workflow for evolving each generation.
3 Experimental Design
3.1 Optimizers & Test Functions
@@
This choice is justified by the fact that the QUASAR algorithm is built on the DE framework, as well as DE’s standing as a robust and effective optimizer [5]. L-SHADE was specifically chosen due to its success in CEC optimization competitions [4].
@@
3.2 Experimental Setup
To ensure a fair comparison, each optimizer was run using their default hyperparameters. The initial population generation methods were also left to defaults; the tested sample sizes not being powers of 2 diminishes the uniformity of QUASAR’s initial Sobol sequence [2], providing a theoretical initial handicap.
@@ 3.2 Experimental Setup
The computational budget for metaheuristic comparison is often defined by the maximum number of objective function evaluations ($FES_{max}$) [3]. Due to L-SHADE’s linear population reduction, the $FES_{max}$ becomes a difficult comparison metric to compare against. As such, the primary comparisons are performed on final solution fitness errors and overall optimization times, given the same number of generations, using modern implementations via SciPy’s DE and MealPy’s L-SHADE.
@@ 3.3 Experimental Trials:
3.3.1 Dimension-Variant Trials:
The dimension-variant trials assessed QUASAR’s scalability by scaling the problem dimension $D$ with a fixed population size ($N=1000$). 3030 independent trials were performed for each test function/dimension combination to ensure statistical significance.
@@ 3.3.1 Dimension-Variant Trials:
3.3.2 Sample-Variant Trials:
The sample-variant trials assessed the impact of population size on performance. The computational budget is varied between population sizes of $[100,250,500,1000,5000]$. 3030 independent trials were performed for each test function / sample size combination, using maximum number of iterations $g_{max}=100$ and fixed dimensionality $D=100$.
@@ 3.3.2 Sample-Variant Trials:
Table 1: Experimental Parameters for Sample and Dimension Variations
@@
3.4 Statistical Analysis:
3.4.1 Overall Rank:
@@
3.4.2 Solution Quality Improvements:
@@ 3.4.3 Computational Efficiency:
$\bar{R}_{\text{time}}=\sqrt[k]{\prod{j=1}^{k}R_{\text{time},j}}$ (11)
3.4.4 Statistical Significance:
@@ 4 Results
Table 2: Rank Sums by Sample Size and Dimension
@@
Table 3: Overall Performance Comparison (Geometric Mean Ratios)
@@ 4.1.1 Dimension-Variant:
Figure 3: Geometric mean error (log-scale) by dimension for each optimizer, with population size of 1000.
4.1.2 Sample-Variant:
To assess the impact of varying population sizes, the optimizers were compared across a fixed $100D$ and 100 maximum generations $g_{max}$, using various sample sizes ($N$).
@@ 4.1.2 Sample-Variant:
Figure 4: Geometric mean error (log-scale) by population size for each optimizer, in 100 dimensions.
Table 4: Solution Quality Comparison (Geometric Mean Error)
@@ 4.2.1 Dimension-Variant Efficiency:
Figure 5: Average optimization run times for each optimizer, by dimension.
4.2.2 Sample-Variant Efficiency:
@@
Figure 6: Average optimization run times for each optimizer, by population size.
Table 5: Efficiency Comparison (Geometric Mean Run Time)
@@
5 Ease of Use and Implementation
5.1 Ease of Use
@@ 5.2 Implementation
The source code requires minimal dependencies, relying only on NumPy and SciPy. The open-source package containing QUASAR, hdim_opt, is available at https://github.com/jgsoltes/hdim-opt, which includes the experimental notebooks for full reproducibility.
6 Conclusion
@@
7 Future Works
QUASAR’s quasi-adaptive mechanisms may be improved via strategies found in other modern DE variants.
7.1 Enhancements
7.1.1 Mutations:
@@ 7.1.4 Verification:
Acknowledgements
The author would like to thank Dr. Kellen Sorauf, Dr. Ksenia Polson, Dr. Mike Busch, and Monet Morris for their invaluable feedback and support throughout this study. It would not have been possible without the insights gained from each of our discussions.
Dr. Sorauf’s recommendation to include L-SHADE in the experiment significantly improved the rigor of the study; Dr. Polson’s lectures laid the framework for the statistical analysis; Dr. Busch pushed me to pursue the project to begin with. Monet remains a source of inspiration, as always.
This research was conducted independently and was not part of a funded project.